# A Representation Transformation of Parametric Feynman Integrals

### v2

@@ V Discussion
The transformation and the corresponding duality can be applied to various aspects of Feynman integrals, including Laundau analysis and integral reduction. Specifically, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation. Additionally, eq. III.5 establishes a correspondence between a syzygy equation in the Baikov representation [57, 58] and a parametric annihilator in the parametric representation [6, 17].
@@ Acknowledgements.
$0=$ $\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}$
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}$
$=$ $(L+1)\Lambda_{i,(n+1)(n+1)}$
$+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\left(\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}\right)u_{j}x_{k}$
$\equiv$ $(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{}.$
@@
$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
with $f(x)$ a polynomial in $x$. Then we have
@@
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)-i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
@@
Thus
@@
$I_{0}(1)=C_{\mathcal{F}}v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},$ (A.1)
where $C_{\mathcal{F}}$ is a constant. The $v_{n+1}$ dependence is fixed through a simple power counting.
Appendix B Derivation of eq. (IV.10)
@@
$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
Thus we have
$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$
@@
$\left(A^{-1}\right)_{ij}=\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.$
@@
$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.1)
@@
$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.2a)
$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$ (B.2b)
×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.2c)
These equations are of the form of the complete condition eq. III.5, since $A_{ij}$ and $B_{ij}$ are linear in $x$. It is easy to see that they contain $\frac{1}{2}L(L+1)+E=n$ independent equations. $\sum_{j=1}^{n}\Lambda_{i,jj}$ can be obtained from eq. III.5 by replaceing $u_{j}x_{k}$ by $\delta_{jk}$. Thus eq. B.2a gives [Remember the definitions in eq. IV.6.]
$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{}.$
And eq. B.2c gives (Notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)

### v3

@@ V Discussion
The transformation and the corresponding duality can be widely applied to the studies of Feynman integrals, including integral reduction and Laundau analysis. Specifically, eq. III.5 establishes a correspondence between a syzygy equation in the Baikov representation [57, 58] and a parametric annihilator in the parametric representation [6, 17]. Additionally, the transformation in II.1 establishes a duality between the first (second) Landau equation in the parametric representation and the second (first) Landau equation in the Baikov representation. It is known that the infrared singularities of Feynman integrals can be identified with solutions of Landau equations. The method of regions [68, 69, 70, 71, 72] provides a systematic method to isolate the infrared structures of Feynman integrals. While endpoint singularities (in the parametric space) can be identified using a geometric method [71, 73, 74], this method can not directly be applied to more general regions, including threshold regions [68, 75, 76, 77]. The latter are related to solutions of the second Landau equations which are away from the origin. By virtue of the duality found in this letter, it is expected that this kind of regions can be identified by applying the geometric method to the dual representation. This is confirmed through some one-loop integrals, including those considered in refs. [68, 75]. The application of this method to multiloop integrals is complicated by the presence of irreducible scalar products, which deserves further investigation.
@@ Acknowledgements.
$\begin{split}0=&\sum_{j,k=1}^{n+1}\Lambda_{i,jk}u_{j}x_{k}\\ =&\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda_{i,jk}u_{j}x_{k}+\Lambda_{i,(n+1)(n+1)}u_{n+1}x_{n+1}\\ =&(L+1)\Lambda_{i,(n+1)(n+1)}+\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}u_{j}x_{k}~{},\end{split}$ (A.1)
with
$\Lambda^{\prime}_{i,jk}\equiv\Lambda_{i,jk}-\Lambda_{i,(n+1)(n+1)}\delta_{jk}~{}.$
@@
$I_{0}(f)\equiv\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\mathcal{F}^{\lambda_{0}}f(x)\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)~{},$
with $f(x)$ a polynomial in $x$. Then, by virtue of eq. (A.1), we have
@@
$=$ $\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}\int_{-\infty}^{\infty}\prod_{i=1}^{n+1}\mathrm{d}x_{i}~{}\frac{\partial\mathcal{F}^{\lambda_{0}}}{\partial x_{j}}x_{k}\exp\left(-i\sum_{i=1}^{n+1}v_{i}x_{i}\right)$
$=$ $-\sum_{j=1}^{n}\Lambda^{\prime}_{i,jj}I_{0}(1)+i\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}I_{0}(x_{k})$
@@
On the other hand, eq. (A.1) is equivalent to
$\sum_{j=1}^{n+1}\sum_{k=1}^{n}\Lambda^{\prime}_{i,jk}v_{j}\frac{\partial\mathcal{G}(v)}{\partial v_{k}}=-(L+1)\Lambda_{i,(n+1)(n+1)}\mathcal{G}~{}.$
Thus we get
@@
$\begin{split}I_{0}(1)=&(-2i\pi)^{n+1}e^{\frac{i}{2}\pi(L+1)\lambda_{0}}C_{\mathcal{F}}\\ &\times v_{n+1}^{-n_{0}-1}\mathcal{G}(v)^{-\lambda_{0}-\frac{n-n_{0}}{L+1}}~{},\end{split}$ (A.2)
where $C_{\mathcal{F}}$ is a constant, and the factors in $i$ and $\pi$ are added for future convenience. The $v_{n+1}$ dependence is fixed through a simple power counting.
Appendix B Derivation of eq. (IV.7)
@@
$=$ $2\left(A^{-1}\right)_{ik}B_{kl}p_{l}\cdot p_{j}u_{n+1}\mathcal{F}$
$-\beta_{ij,k}b_{kl}C_{l}u_{n+1}\mathcal{F}~{}.$
Multiplying both sides with $\frac{1}{2u_{n+1}\mathcal{F}}A_{ij}g_{kl}$, we get
$B_{ij}=A_{ik}g_{jl}B^{\prime}_{kl}~{}.$ (B.1)
@@
$\begin{split}\left(A^{-1}\right)_{ij}=&\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\\ &+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}~{}.\end{split}$ (B.2)
@@
$\begin{split}\mathcal{G}=&\mathcal{F}^{-1}=u_{n+1}U^{-1}=u_{n+1}\det\left(A^{-1}\right)\\ =&u_{n+1}\det\left[\alpha_{ij,k}a_{kl}(u_{l}+u_{n+1}C_{l})\right.\\ &\left.+u_{n+1}g_{mn}B^{\prime}_{im}B^{\prime}_{jn}\right]~{}.\end{split}$ (B.3)
@@
$0=$ $A_{ik}\beta_{kj,l}b_{lm}(u_{m}+C_{m}u_{n+1})$
$-2B_{ik}p_{k}\cdot p_{j}u_{n+1}~{},$ (B.4a)
$0=$ $(L+1)\left(A_{ik}\alpha_{kj,l}a_{lm}+\frac{1}{2}B_{ik}\beta_{jk,l}b_{lm}\right)$
×$\times(u_{m}+C_{m}u_{n+1})-\delta_{ij}u_{k}x_{k}~{}.$ (B.4b)
Here, the first equation is obtained from eq. (B.1) by multiplying both sides with $u_{n+1}$, and the second equation is obtained from eq. (B.2) by multiplying both sides with the matrix $A_{ij}$ and replacing $\delta_{ij}$ by $\frac{1}{L+1}\delta_{ij}u_{k}x_{k}$. These equations are of the form of the complete condition eq. III.5, since $A_{ij}$ and $B_{ij}$ are linear in $x$. It is easy to see that they contain $\frac{1}{2}L(L+1)+LE=n$ independent equations. $\sum_{j=1}^{n}\Lambda_{i,jj}$ can be obtained from eq. III.5 by replacing $u_{j}x_{k}$ by $\delta_{jk}$. Thus eq. B.4a gives [remember the definitions in eqs. (IV.6).]
$\sum_{j=1}^{n}\Lambda_{i,jj}=\Lambda_{i,(n+1)(n+1)}=0~{},$
and eq. B.4b gives (notice that the matrix $A_{ij}$ is of dimension $L\times L$, and the matrix $B_{ij}$ is of dimension $L\times E$.)